# MetaToken: Detecting Hallucination in Image Descriptions by Meta Classification

### v1

@@ 3.2 Input Metrics
    $A_{o_{j}}^{g}=\frac{1}{u+1}\sum_{r=0}^{u}|Attention_{t_{o_{j,s}}}({t_{x_{r}}})|,$ (5)
    where $Attention_{t_{i}}({t_{n}})$ denotes the attention of the generated token $t_{i}$ at generation step $i$ on the input token $t_{n}$,
@@
  • the sequence score [^1] https://huggingface.co/docs/transformers/main/en/main_classes/text_generation#transformers.GenerationMixin.compute_transition_scores with length_penalty parameter $l_{p}$
@@
    $M_{o_{j}}=1-p_{o_{j,s}}(t_{\mathrm{max}})+\max_{t\in\mathcal{V}\setminus\{t_{\mathrm{max}}\}}p_{o_{j,s}}(t),$ (12)
@@
    $D_{o_{j}}=\log p_{o_{j,s}}(t_{\mathrm{max}})-\log\hat{p}_{o_{j,s}},\quad\textrm{and},$ (13)
  • for LVLMs which use the same vision encoder as CLIP [50], the CLIPScore [21]
    $CLIP_{o_{j}}=\max(100*\cos(\mathcal{E}_{x},\mathcal{E}_{{Cap}_{o_{j}}}),0)$ (14)
    with the visual CLIP embedding $\mathcal{E}_{x}$ for image $x$ and the textual CLIP embedding $\mathcal{E}_{{Cap}_{o_{j}}}$ for the caption ${Cap}_{o_{j}}$ which is defined as "a photo of a <$o_{j}$>".
Finally, for an object $o_{j}\in\mathcal{O}_{s}$, we define the total set of input metrics for general LVLMs as
$\begin{split}\mathcal{M}_{o_{j}}=\{I_{o_{j}},H_{o_{j}},P_{o_{j}},N_{o_{j}},A^{0}_{o_{j}},\dots,A^{G-1}_{o_{j}},\\ L_{o_{j}},C_{o_{j}},S_{o_{j}},V_{o_{j}},E_{o_{j}},R_{o_{j}},M_{o_{j}},D_{o_{j}}\}\end{split}$ (15)
with cardinality $|\mathcal{M}_{o_{j}}|=12+G$ and for LVLMs that share the visual embedding with CLIP, the extended set of metrics as
$\begin{split}\mathcal{M}^{\mathrm{clip}}_{o_{j}}=\{I_{o_{j}},H_{o_{j}},P_{o_{j}},N_{o_{j}},A^{0}_{o_{j}},\dots,A^{G-1}_{o_{j}},\\ L_{o_{j}},C_{o_{j}},S_{o_{j}},V_{o_{j}},E_{o_{j}},R_{o_{j}},M_{o_{j}},D_{o_{j}},CLIP_{o_{j}}\}\end{split}$ (16)
with cardinality $|\mathcal{M}^{\mathrm{clip}}_{o_{j}}|=13+G$.
3.3 Hallucination Detection
In this section, we assume a general LVLM that uses a vision encoder different from CLIP. In order to detect hallucinations in the generated output $s$, we train a lightweight binary meta model to classify between true and hallucinated objects. To this end, let
$f:\mathbb{R}^{12+G}\rightarrow\{0,1\}$ (17)
denote the binary classifier. We denote our set of training captions by $\mathcal{S}^{\mathrm{train}}$ and the corresponding set of generated objects by
@@ 3.3 Hallucination Detection
For every object $o_{j}\in\mathcal{O}_{\mathcal{S}^{\mathrm{train}}}$ we build an input vector $m_{o_{j}}\in\mathbb{R}^{12+G}$ representing the metric set $\mathcal{M}_{o_{j}}$ (Eq. 15) and define the label $y_{o_{j}}\in\{0,1\}$ according to the CHAIR evaluation (see Sec. 4.2). After standardizing the inputs, we use the set
@@
Given a generated caption $s$ at inference time, we calculate the input vector $m_{o_{j}}\in\mathbb{R}^{12+G}$ for every object $o_{j}\in\mathcal{O}_{s}$ and apply the trained binary meta classifier $f$ to detect hallucinated objects. Note that the input vector $m_{o_{j}}$ can be calculated based on the model output only without any knowledge of the ground truth data.
@@ 4 Experimental Settings
Table 2: Generation Configurations. The generation configurations applied in our experiments for nucleus sampling [22] and beam search [3], respectively.
parameter nucleus sampling beam search
do_sample True False
top_p 0.9 0.9
temperature 1 1
num_beams 1 5
max_length 256 256
min_length 1 1
repetition_penalty 1 1.5
length_penalty 1 1
4.1 Dataset
The MSCOCO dataset [37] is a large-scale dataset for object detection, segmentation, and image captioning comprising more than $200K$ labeled images for $80$ object categories. Following [36], we randomly sample $5,000$ images from the MSCOCO validation set and produce image captions $s$ for four state-of-the-art LVLMs. We use $80\%$ of the generated captions as our training set $\mathcal{O}_{\mathcal{S}^{\mathrm{train}}}$ and validate our method on the remaining $20\%$ denoted as $\mathcal{O}_{\mathcal{S}^{\mathrm{val}}}$. In our experiments, we average our results over ten randomly sampled training-validation splits. In tables, the corresponding standard deviations are given in parentheses.
@@ 4.2 Hallucination Evaluation
Given an image caption, the CHAIR method [52] provides a binary label for every generated MSCOCO object and corresponding synonyms [44] indicating whether the object is true, i.e., contained in the image, or hallucinated. The CHAIR metric measures the proportion of hallucinated MSCOCO objects by matching the generated text against the ground truth objects provided in the MSCOCO image captioning and object detection datasets [37]. It is defined on two semantic levels: The object instance level $\textrm{CHAIR}_{i}$ to measure the proportion of hallucinated MSCOCO objects in an image caption which is defined as
Since we tackle the problem of object instance hallucinations in generated captions $s$, we use the $\textrm{CHAIR}_{i}$ metric to extract all MSCOCO objects $\mathcal{O}_{s}=\{o_{0},\dots,o_{z}\}$ and simultaneously label the extracted objects as either true, which is encoded as 00, or hallucinated, which is encoded as $1$.
4.3 Large Vision Language Models
We evaluate our approach on four state-of-the-art open-source LVLMs, i.e., InstructBLIP (Vicuna-7B) [8], mPLUG-Owl (LLaMA-7B) [67], MiniGPT-4 (Vicuna-7B) [73], and LLaVa 1.5 (Vicuna-7B) [23], all of them using $G=32$ attention heads. We use nucleus sampling [22] and the prompt
"Describe all objects in the image."
for all image caption generations. For InstructBLIP and mPLUG-Owl, we create further image captions based on beam search [3]. Tab. 2 states the detailed generation configurations of our experiments. The performance of the LVLMs considered with respect to the average number of generated MSCOCO objects per image as well as the hallucination rate in terms of $\textrm{CHAIR}_{i}$ (Eq. 20) and $\textrm{CHAIR}_{s}$ (Eq. 21) is summarized in Tab. 3. Note that these results highly depend on the respective generation configurations, the underlying LLM (7B vs. 13B), as well as the chosen prompt [36, 62].
[…]
InstructBLIP (sample) $5.6$ 10.4 $33.6$
mPLUG-Owl (sample) $6.9$ $30.2$ $71.6$
MiniGPT-4 (sample) $4.4$ $13.6$ 33.4
LLaVa (sample) $7.3$ $18.7$ $59.2$
InstructBLIP (beam) $5.1$ 7.6 25.0
mPLUG-Owl (beam) $6.7$ $25.3$ $63.4$
Table 4: Experimental Results based on Nucleus Sampling [22]. Hallucination detection results on four state-of-the-art LVLMs with log probability baseline $L$ (Eq. 6) and entropy baseline $E$ (Eq. 10). Ours refers to the metric set $\mathcal{M}$ (Eq. 15), while $\textrm{Ours}^{\textrm{clip}}$ is based on the extended metric set $\mathcal{M}^{\textrm{clip}}$ (Eq. 16). The classifiers based on logistic regression and gradient boosting are denoted by LR and GB, respectively. The best results in each block are highlighted.
LR GB LR GB LR GB
InstructBLIP $L$ $89.46^{(\pm 1.4\mathrm{e}-3)}$ $89.46^{(\pm 1.3\mathrm{e}-3)}$ $73.51^{(\pm 8.7\mathrm{e}-3)}$ $73.16^{(\pm 9.0\mathrm{e}-3)}$ $27.07^{(\pm 2.1\mathrm{e}-2)}$ $25.6^{(\pm 2.5\mathrm{e}-2)}$
$E$ $89.49^{(\pm 1.6\mathrm{e}-3)}$ $89.48^{(\pm 1.6\mathrm{e}-3)}$ $65.49^{(\pm 1.3\mathrm{e}-2)}$ $66.23^{(\pm 1.5\mathrm{e}-2)}$ $15.38^{(\pm 5.7\mathrm{e}-3)}$ $17.68^{(\pm 7.0\mathrm{e}-3)}$
Ours $91.33^{(\pm 1.8\mathrm{e}-3)}$ $\textbf{91.48}^{(\pm 1.5\mathrm{e}-3)}$ $89.94^{(\pm 9.1\mathrm{e}-3)}$ $\textbf{90.39}^{(\pm 6.8\mathrm{e}-3)}$ $56.13^{(\pm 1.2\mathrm{e}-2)}$ $\textbf{57.50}^{(\pm 1.0\mathrm{e}-2)}$
mPLUG-Owl $L$ $72.42^{(\pm 4.3\mathrm{e}-3)}$ $72.48^{(\pm 4.6\mathrm{e}-3)}$ $71.75^{(\pm 9.4\mathrm{e}-3)}$ $71.86^{(\pm 9.3\mathrm{e}-3)}$ $51.21^{(\pm 1.2\mathrm{e}-2)}$ $50.65^{(\pm 1.1\mathrm{e}-2)}$
$E$ $70.06^{(\pm 4.9\mathrm{e}-3)}$ $70.77^{(\pm 2.9\mathrm{e}-3)}$ $66.01^{(\pm 6.3\mathrm{e}-3)}$ $68.33^{(\pm 6.1\mathrm{e}-3)}$ $40.09^{(\pm 8.2\mathrm{e}-3)}$ $45.54^{(\pm 1.2\mathrm{e}-2)}$
Ours $82.87^{(\pm 2.7\mathrm{e}-3)}$ $\textbf{83.56}^{(\pm 3.1\mathrm{e}-3)}$ $88.55^{(\pm 3.7\mathrm{e}-3)}$ $\textbf{89.87}^{(\pm 2.3\mathrm{e}-3)}$ $76.11^{(\pm 5.8\mathrm{e}-3)}$ $\textbf{78.61}^{(\pm 7.9\mathrm{e}-3)}$
$\textrm{Ours}^{\mathrm{clip}}$ $85.05^{(\pm 2.8\mathrm{e}-3)}$ $\textbf{85.84}^{(\pm 2.9\mathrm{e}-3)}$ $91.29^{(\pm 3.2\mathrm{e}-3)}$ $\textbf{92.19}^{(\pm 2.1\mathrm{e}-3)}$ $80.98^{(\pm 4.3\mathrm{e}-3)}$ $\textbf{82.69}^{(\pm 6.1\mathrm{e}-3)}$
MiniGPT-4 $L$ $86.91^{(\pm 3.6\mathrm{e}-3)}$ $86.85^{(\pm 3.9\mathrm{e}-3)}$ $67.26^{(\pm 2.1\mathrm{e}-2)}$ $67.01^{(\pm 2.1\mathrm{e}-2)}$ $26.25^{(\pm 1.7\mathrm{e}-2)}$ $25.41^{(\pm 1.2\mathrm{e}-2)}$
$E$ $86.84^{(\pm 3.6\mathrm{e}-3)}$ $86.82^{(\pm 3.6\mathrm{e}-3)}$ $60.78^{(\pm 1.8\mathrm{e}-2)}$ $63.19^{(\pm 1.2\mathrm{e}-2)}$ $15.77^{(\pm 6.7\mathrm{e}-3)}$ $18.98^{(\pm 1.3\mathrm{e}-2)}$
Ours $88.92^{(\pm 3.5\mathrm{e}-3)}$ $\textbf{89.27}^{(\pm 4.9\mathrm{e}-3)}$ $88.16^{(\pm 1.5\mathrm{e}-2)}$ $\textbf{89.74}^{(\pm 1.3\mathrm{e}-2)}$ $54.90^{(\pm 6.5\mathrm{e}-2)}$ $\textbf{57.27}^{(\pm 5.7\mathrm{e}-2)}$
LLaVa $L$ $81.57^{(\pm 1.4\mathrm{e}-3)}$ $81.49^{(\pm 1.6\mathrm{e}-3)}$ $70.53^{(\pm 8.7\mathrm{e}-3)}$ $70.73^{(\pm 6.6\mathrm{e}-3)}$ $37.53^{(\pm 2.0\mathrm{e}-2)}$ $36.59^{(\pm 1.7\mathrm{e}-2)}$
$E$ $81.28^{(\pm 2.8\mathrm{e}-3)}$ $81.26^{(\pm 2.9\mathrm{e}-3)}$ $62.73^{(\pm 9.0\mathrm{e}-3)}$ $64.63^{(\pm 7.7\mathrm{e}-3)}$ $23.85^{(\pm 6.3\mathrm{e}-3)}$ $27.52^{(\pm 4.6\mathrm{e}-3)}$
Ours $87.25^{(\pm 2.0\mathrm{e}-3)}$ $\textbf{87.78}^{(\pm 3.0\mathrm{e}-3)}$ $90.05^{(\pm 4.0\mathrm{e}-3)}$ $\textbf{91.01}^{(\pm 4.3\mathrm{e}-3)}$ $70.15^{(\pm 1.0\mathrm{e}-2)}$ $\textbf{72.58}^{(\pm 1.3\mathrm{e}-2)}$
$\textrm{Ours}^{\mathrm{clip}}$ $87.93^{(\pm 1.2\mathrm{e}-3)}$ $\textbf{88.36}^{(\pm 1.6\mathrm{e}-3)}$ $91.73^{(\pm 2.4\mathrm{e}-3)}$ $\textbf{92.50}^{(\pm 2.5\mathrm{e}-3)}$ $73.00^{(\pm 7.8\mathrm{e}-3)}$ $\textbf{75.08}^{(\pm 8.5\mathrm{e}-3)}$
4.4 Evaluation Metrics and Meta Models
Table 5: Expected Calibration Error. The $ECE$ for the logistic regression (LR) and gradient boosting (GB) MetaToken classifier.
@@ 4.4 Evaluation Metrics and Meta Models
We evaluate our method based on the accuracy $ACC$, the area under receiver operator characteristic curve $AUROC$ [10] and the area under precision recall curve $AUPRC$ [10]. The receiver operator characteristic curve illustrates the performance of a binary classifier by plotting the true positive rate against the false positive rate at various decision thresholds indicating the ability to distinguish between both classes. The precision recall curve plots precision values against the recall at various decision thresholds accounting for imbalance in the underlying dataset. Since we observe highly imbalanced data with respect to object instance hallucinations (see Tab. 3), the main focus in our evaluation is on the $AUROC$ and $AUPRC$ value. We compare two binary meta models, i.e., a classifier based on a logistic regression and a meta model based on gradient boosting.
@@ 4.5 Baseline
As previous works on meta classification [47, 53, 54], we use common uncertainty measures as our baseline. In our experiments, we consider the log probability baseline $L$ (Eq. 6) and the entropy baseline $E$ (Eq. 10). For both metrics, we train a baseline classifier in the one-dimensional space, i.e.,
@@
with training set $\{(L_{o_{j}},y_{o_{j}})\quad|\quad j=0,\dots,|\mathcal{O}_{\mathcal{S}^{\textrm{train}}}|-1\}$ and $\{(E_{o_{j}},y_{o_{j}})\quad|\quad j=0,\dots,|\mathcal{O}_{\mathcal{S}^{\textrm{train}}}|-1\}$, respectively.
Note that a direct comparison of our approach to the state-of-the-art methods listed in Tab. 1 is not possible. While our method tackles the problem of token-level instance hallucination, the listed methods either evaluate image captions on a sentence- or subsentence-level [19, 62] or based on atomic facts extracted from the image captions [65, 68, 26, 6]. Thus, the proposed methods [19] and [62] do not provide any information on which specific word of the respective sentence or subsentence is hallucinated. Similarly, [65, 68, 26, 6] are based on atomic facts which neither allow for a token-level evaluation. Moreover, the analysis in [26] shows that the atomic fact extraction already induces detection errors propagating through the detection and evaluation pipeline. To overcome these issues, we rely on the handcrafted CHAIR evaluation (see Sec. 4.2) based on the human-labeled MSCOCO dataset.
Table 6: Experimental Results based on Beam Search [3]. Hallucination detection results for InstructBLIP [8] and mPLUG-Owl [67] with log probability baseline $L$ (Eq. 6) and entropy baseline $E$ (Eq. 10). Ours refers to the metric set $\mathcal{M}$ (Eq. 15). The classifier based on logistic regression and gradient boosting are denoted by LR and GB, respectively. The best results in each block are highlighted.

### v2

@@ 3.2 Input Features
    $A_{o_{j}}^{g}=\frac{1}{u+1}\sum_{r=0}^{u}|\textrm{Attention}_{t_{o_{j,s}}}({t_{x_{r}}})|,$ (3)
    where $\textrm{Attention}_{t_{i}}({t_{n}})$ denotes the attention of the generated token $t_{i}$ at generation step $i$ on the input token $t_{n}$,
@@
  • the sequence score with length_penalty parameter $l_{p}$
@@
    $M_{o_{j}}=R_{o_{j}}+\max_{t\in\mathcal{V}\setminus\{t_{\mathrm{max}}\}}p_{o_{j,s}}(t),\quad\textrm{and},$ (10)
@@
    $D_{o_{j}}=\log p_{o_{j,s}}(t_{\mathrm{max}})-\log\hat{p}_{o_{j,s}}.$ (11)
Finally, for an object $o_{j}\in\mathcal{O}_{s}$, we define the total set of input features as
$\begin{split}\mathcal{M}_{o_{j}}=\{&P_{o_{j}},N_{o_{j}},A^{0}_{o_{j}},\dots,A^{G-1}_{o_{j}},L_{o_{j}},\\ &C_{o_{j}},S_{o_{j}},V_{o_{j}},E_{o_{j}},R_{o_{j}},M_{o_{j}},D_{o_{j}}\}\end{split}$ (12)
with cardinality $|\mathcal{M}_{o_{j}}|=10+G$.
3.3 Hallucination Detection Using     Meta Classification
Let $\mathcal{M}$ denote a set of input features with cardinality $|\mathcal{M}|$. The idea of meta classification consists of training a lightweight binary meta model based on the input features $\mathcal{M}$ to classify between true and false predictions, i.e., to detect true and hallucinated objects in the generated output $s$. To this end, let
$f:\mathbb{R}^{|\mathcal{M}|}\rightarrow\{0,1\}$ (13)
denote the binary classifier. We denote the set of training captions by $\mathcal{S}^{\mathrm{train}}$ and the corresponding set of generated objects by
@@ 3.3 Hallucination Detection Using     Meta Classification
For every generated $o_{j}\in\mathcal{O}_{\mathcal{S}^{\mathrm{train}}}$ we build an input vector $m_{o_{j}}\in\mathbb{R}^{|\mathcal{M}_{o_{j}}|}$ representing the feature set $\mathcal{M}_{o_{j}}$ (Eq. 12) and define the label $y_{o_{j}}\in\{0,1\}$ according to the CHAIR evaluation (see Sec. 4.2). After standardizing the inputs, we use the set
@@
Given a generated caption $s$ at inference time, we calculate the input vector $m_{o_{j}}\in\mathbb{R}^{|\mathcal{M}_{o_{j}}|}$ for every object $o_{j}\in\mathcal{O}_{s}$ and apply the trained binary meta classifier $f$ to detect hallucinated objects. Note that the input vector $m_{o_{j}}$ can be calculated in an automated manner based on the model output only without any knowledge of the ground truth data.
[…]
MSCOCO BDD100K
[…]
InstructBLIP (s) $5.6$ 10.4 $5.8$ 22.07
mPLUG-Owl (s) $6.9$ $30.2$ $5.8$ $33.44$
MiniGPT-4 (s) $4.4$ $13.6$ $3.7$ $29.12$
LLaVa (s) $7.3$ $18.7$ $7.1$ $29.72$
InstructBLIP (b) $5.1$ 7.6 - -
mPLUG-Owl (b) $6.7$ $25.3$
@@ 4 EXPERIMENTAL SETTINGS
4.1 Datasets
We evaluate our method on the MSCOCO and BDD100K datasets. The MSCOCO dataset [Lin et al., 2014] is a large-scale dataset for object detection, segmentation, and image captioning comprising more than $200K$ labeled images. The BDD100K dataset [Yu et al., 2020] consists of $100K$ labeled street scene images including labels for object detection, semantic segmentation and instance segmentation. We randomly sample $5,000$ images from the validation sets and produce image captions $s$ for four SOTA LVLMs. We use $80\%$ of the generated captions as training set $\mathcal{O}_{\mathcal{S}^{\mathrm{train}}}$ and validate our method on the remaining $20\%$ denoted as $\mathcal{O}_{\mathcal{S}^{\mathrm{val}}}$. In our experiments, we average our results over ten randomly sampled training-validation splits. The corresponding standard deviations are given in parentheses.
@@ 4.2 Hallucination Evaluation
Table 3: Expected Calibration Error. The $ECE$ for the LR and GB MetaToken classifier.
@@
The CHAIR metric [Rohrbach et al., 2018] is an automated hallucination evaluation method which measures the proportion of hallucinated objects by matching the generated text against the ground truth objects. Given an image caption $s$, CHAIR provides a binary label for the generated objects $o_{j}\in\mathcal{O}_{s}$ indicating whether each object is true, i.e., contained in the image, or hallucinated. CHAIR is defined on two semantic levels: The object instance level $\textrm{CHAIR}_{i}$ to measure the proportion of hallucinated objects in an image caption which is defined as
Since we tackle the problem of object hallucinations in generated captions $s$, we use the $\textrm{CHAIR}_{i}$ metric to extract the objects $\mathcal{O}_{s}=\{o_{0},\dots,o_{z}\}$ and simultaneously label the extracted objects as either true, which is encoded as 00, or hallucinated, which is encoded as $1$. For more details, see Sec. 8.2.
4.3 Large Vision Language Models
We evaluate our approach on four SOTA open-source LVLMs, i.e., InstructBLIP (Vicuna-7B) [Dai et al., 2023a], mPLUG-Owl (LLaMA-7B) [Ye et al., 2023], MiniGPT-4 (Vicuna-7B) [Zhu et al., 2023], and LLaVa 1.5 (Vicuna-7B) [Huang et al., 2023], all of them using $G=32$ attention heads. We use nucleus sampling [Holtzman et al., 2020] and the prompt
"Describe all objects in the image."
for all image caption generations. For InstructBLIP and mPLUG-Owl, we create further image captions on MSCOCO using beam search [Bisiani, 1987]. Tab. 9 states the detailed generation configurations of our experiments. The performance of the LVLMs considered with respect to the average number of generated objects per image as well as the hallucination rate in terms of $\textrm{CHAIR}_{i}$ (Eq. 16) is summarized in Tab. 2. Note that these results highly depend on the respective generation configurations, the underlying LLM (7B vs. 13B), as well as the chosen prompt [Li et al., 2023b, Wang et al., 2023].
4.4 Evaluation Metrics and           Meta Models
We evaluate our method based on the accuracy $ACC$, the area under receiver operator characteristic curve $AUROC$ [Davis and Goadrich, 2006] and the area under precision recall curve $AUPRC$ [Davis and Goadrich, 2006]. The receiver operator characteristic curve illustrates the performance of a binary classifier by plotting the true positive rate against the false positive rate at various decision thresholds indicating the ability to distinguish between both classes. The precision recall curve plots precision values against the recall at various decision thresholds accounting for imbalance in the underlying dataset. Since we observe imbalanced data with respect to object instance hallucinations (see Tab. 2), the main focus in our evaluation is on the $AUROC$ and $AUPRC$ value. We compare two binary meta models, i.e., a classifier based on a logistic regression (LR) and a gradient boosting (GB) meta model.
@@ 4.5 Baseline
We use the reference-free token-level algebraic confidence TLC-A [Petryk et al., 2023] as our baseline. We consider the log probability-based token-level confidence $L$ (see Eq. 4) and the entropy-based confidence $E$ (see Eq. 8). For both confidence measures, we train a baseline classifier in the one-dimensional space, i.e.,
@@
with training set $\{(L_{o_{j}},y_{o_{j}})|j=0,\dots,|\mathcal{O}_{\mathcal{S}^{\textrm{train}}}|-1\}$ and $\{(E_{o_{j}},y_{o_{j}})|j=0,\dots,|\mathcal{O}_{\mathcal{S}^{\textrm{train}}}|-1\}$, respectively.
Note that a direct comparison of our approach to the detection methods listed in Tab. 1 is not possible. While our method tackles the problem of token-level object hallucination, the listed methods either evaluate hallucinations on a sentence- or subsentence-level on their own human-labeled dataset [Gunjal et al., 2023, Wang et al., 2023] or based on atomic facts extracted from the image captions using LLMs, which are then labeled using LVLMs [Wu et al., 2024, Yin et al., 2023, Jing et al., 2023, Chen et al., 2024]. Thus, the proposed methods [Gunjal et al., 2023] and [Wang et al., 2023] do not provide any information on which specific word of the respective sentence or subsentence is hallucinated. Similarly, [Wu et al., 2024, Yin et al., 2023, Jing et al., 2023, Chen et al., 2024] are based on LLM-generated atomic facts which neither allow for a token-level evaluation and are not reproducible. Moreover, the analysis in [Jing et al., 2023] shows that the LLM-based atomic fact extraction already induces errors propagating through the detection and evaluation pipeline. To overcome these issues, we rely on the automated and reproducible CHAIR evaluation method (see Sec. 4.2).
Table 4: Computational Time. The average time for feature calculation per image (feature), classifier training using $4.000$ image captions (train) and inference on $1.000$ image captions (predict).
feature (sec.) train (sec.) predict (sec.)
LR $0.07174$ $0.47187$ $0.00198$
GB $0.07174$ $54.00010$ $0.00385$
Table 5: Experimental Results based on Nucleus Sampling [Holtzman et al., 2020]. Hallucination detection results on four SOTA LVLMs. Ours refers to the feature set $\mathcal{M}$. The best results in each block are highlighted.
MSCOCO [Lin et al., 2014]